OpenStudy (anonymous):

I will give medal and fan!!! 4 over 12 = x over 18

OpenStudy (anonymous):

@shaik0124

OpenStudy (secret-ninja):

\[\frac{ 4 }{ 12 } = \frac{ x }{ 18 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is it

Nnesha (nnesha):

can't you not cross multiply

Nnesha (nnesha):

what is 4 times 18 ??? what is 12 times x = ??

OpenStudy (anonymous):

ik, i want to know what she got

OpenStudy (secret-ninja):

http://www.wikihow.com/Cross-Multiply

Nnesha (nnesha):

okay first tell me what ou got

OpenStudy (shaik0124):

4/12=x/18 now solve like previous problems

OpenStudy (shaik0124):

i cant say the answer everytime u should learn the ethod ottherwise no use of study

OpenStudy (shaik0124):

nisha said to stop saying answers so i stopped now

Nnesha (nnesha):

its on the coc http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

4 times 18 = 72

OpenStudy (shaik0124):

i knnow

OpenStudy (anonymous):

then the other is 12x

OpenStudy (shaik0124):

4/12= x/18 now first eliminate the denominator of any oneside first

OpenStudy (anonymous):

ok

OpenStudy (shaik0124):

what did u get now

OpenStudy (shaik0124):

@Nnesha is it ook :)

OpenStudy (shaik0124):

:)

Nnesha (nnesha):

 now 12 is multiplying with x you have to do opposite to solve for x divide both side by 12